OpenStudy (anonymous):

which is larger, 42 or 256 Questoins A.42 B. They Equal C.256

OpenStudy (aripotta):

you don't know which is the bigger number?

hartnn (hartnn):

which do u think is larger ?

hartnn (hartnn):

a 3 digit number is always larger than a 2-digit number

hartnn (hartnn):

so @lovevbird83100 which one is larger ? bdw \(\huge \color{red}{\text{Welcome to Open Study}}\ddot\smile\)

OpenStudy (anonymous):

\[^{?} \sqrt{?}256 or 42\]

hartnn (hartnn):

is that \(\huge \sqrt{256}\) ??

OpenStudy (anonymous):

yess

hartnn (hartnn):

do u know 256 is square of which number ?

OpenStudy (anonymous):

42

hartnn (hartnn):

nopes. 16*16 = 256 so \(\sqrt{256}=16\) ok?

hartnn (hartnn):

so which is larger ? 16 or 42 ?